OpenStudy (anonymous):

1. From 9x-2y subtract x+3y 2. Multiply x+4 and 5x-2 3. Factor 7x^2-22x+3

OpenStudy (princeharryyy):

1) 8x-5y 2) (x+4)(5x-2) =5x^2 -2x + 20x -8 =5x^2 + 18x -8 ur third part is not clear i'm not able to understand wat u have written...

OpenStudy (princeharryyy):

@PJtheVet

OpenStudy (anonymous):

3. is 7x squared minus 22x plus 3

OpenStudy (anonymous):

or 7x to the second - 22x +3

OpenStudy (anonymous):

but factoring it

OpenStudy (princeharryyy):

ohk @PJtheVet

OpenStudy (princeharryyy):

7x^2-22x+3 = 7*(x-3)*(x-1/7) @PJtheVet n sorry for getting late... was talking to some1....